According to one survey in India, 75% of Instagram users love REELS. Suppose that 25


Instagram users (randomly selected) have been approached in the university located in vile


parle. They have been asked about their status of like/ dislike the Instagram- REELS.


a) What is the probability that Exactly 15 of them would agree with the claim (or said they


love Insta-REELS)?


b) What is the probability that Exactly 20 of them would agree with the claim (or said they


love Insta-REELS)?

1
Expert's answer
2022-05-18T06:40:30-0400

Let "X=" the number of  users who love REELS: "X\sim Bin(n, p)."

Given "n=205, p=0.75, q=1-p=0.25."

a)


"P(X=15)=\dbinom{25}{15}(0.75)^{15}(0.25)^{25-15}""=0.04165835"

b)


"P(X=20)=\dbinom{25}{20}(0.75)^{20}(0.25)^{25-20}""=0.16453759"